Two identical cells each of e. m. f E and internal resistance r are first connected in series, and next in parallel. in each case they are applied across the two ends of a wire resistance R. it is found that the same current flows through the wire. what must be the relation between R and r?
1
Expert's answer
2019-04-02T10:44:56-0400

In case of a serial connection of two identical e.m.f the current flowing in the circuit is


"I_1 = \frac{2E}{2r + R}"

In case of a parallel connection of two identical e.m.f. the current is


"I_2 = \frac{E}{\frac{r}{2} + R}"

Following the condition I1 = I2, we obtain


"\frac{2E}{2r + R} = \frac{E}{\frac{r}{2} + R}"

Solving this equation in respect to r, one can derive:


"r = R"

Answer: both r and R should be of the same value.